OpenStudy (anonymous):

Assume that lim x->c = [f(x)+g(x)] = 2, lim x->c [f(x)-g(x)] = 1and that lim x->c f(x) and lim x->c g(x) exsist. find those limits?

OpenStudy (anonymous):

f(c) + g(c)=2...............i f(c) - g(c)=-1.............ii

OpenStudy (anonymous):

Now, solve eqn i and ii to get f(c) and g(c)

OpenStudy (anonymous):

DOES IT HELP?

OpenStudy (anonymous):

opps f(c) - g(c)=1.............ii

OpenStudy (anonymous):

ur a goddam retard

OpenStudy (anonymous):

Your head is broken dumbass

OpenStudy (anonymous):

u is stupid

OpenStudy (anonymous):

???

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

what are all those comments up there? ^

OpenStudy (anonymous):

it isn't making sense? how will I solve?

OpenStudy (anonymous):

\[\lim_{x\to c}[f(x)+g(x)] = 2\\\lim_{x\to c}[f(x)-g(x)] = 1\] \[\lim_{x\to c}[f(x)+g(x)] = 2\\\lim_{x\to c}f(x)+\lim_{x\to c}[g(x)] = 2\] \[\lim_{x\to c}[f(x)-g(x)] = 1\\\lim_{x\to c}f(x)-\lim_{x\to c}[g(x)] = 1\] \[\lim_{x\to c}f(x)+\lim_{x\to c}[g(x)] = 2\\\lim_{x\to c}f(x)-\lim_{x\to c}[g(x)] = 1\\2\lim_{x\to c}f(x)=3\\\lim_{x\to c}f(x)=\frac32\] \[\lim_{x\to c}f(x)-\lim_{x\to c}[g(x)] = 1\\\lim_{x\to c}g(x)=\frac32-1=\frac12\]